OpenStudy (anonymous):

Please Solve 4^x = 256 And if you can please show your work so I can learn how to do it.

OpenStudy (anonymous):

4^4 = 256 = 4^x, so x =4

OpenStudy (ja1):

Um that does not seem right

OpenStudy (anonymous):

Its correct dude.Plz check once..!!

OpenStudy (ja1):

Yah that is correct, sorry I didn't pay attention to your reasoning :)

OpenStudy (anonymous):

No problem dude..!! U r always welcome..:-)

OpenStudy (ja1):

Yah I took this as a different question :)